OpenStudy (anonymous):

I have no idea how to solve this. Can someone please help me for a medal? :) (posting below)

OpenStudy (anonymous):

\[\frac{ f(x+Δx)−f(x) }{ Δx }\] if f(x) = x^2

OpenStudy (irishboy123):

compute as \(\huge \frac{(x + \Delta x)^2 - x^2}{\Delta x}\)

OpenStudy (irishboy123):

so expand out the \(\large (x + \Delta x)^2\) in the numerator :p

OpenStudy (anonymous):

I know how to get to the expanding part, but I am not sure how to expand the numerator. Could you show me how?

OpenStudy (anonymous):

Would it end up looking like this... \[\frac{ x^{2} + Δx^{2}-x^{2} }{ Δx } \]

OpenStudy (irishboy123):

first expand out \( (x+Δx)^2\) on its own that's \( (x+Δx) \times (x+Δx)\)

OpenStudy (irishboy123):